OpenStudy (anonymous):

A system of equations is shown below: 6x – 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x – 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x – y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@k_lynn , @CGGURUMANJUNATH , @ganeshie8

OpenStudy (anonymous):

I would say the last option satisfies the question once it is worked out x =17/28

OpenStudy (anonymous):

The last one?

OpenStudy (anonymous):

@Lena772